OpenStudy (chazz16):

Help MEDAL WILL BE AWARDED!!!! 3^9 * 3^-3

OpenStudy (igreen):

When multiplying like bases with exponents, add the exponents and keep the base. \(\sf x^a \cdot x^b \rightarrow x^{a+b}\)

OpenStudy (igreen):

So in this case, we have: \(\sf 3^{~9~+~(-3)}\)

OpenStudy (igreen):

Can you simplify that?

OpenStudy (anonymous):

729 3^9 = 19,683 (3*3*3*3*3*3*3*3*3) 3^-3 = 1/27 (1/3 * 3 *3, aka 1/27) Then multiply 19,683 by 1/27 Which would be 729.

OpenStudy (anonymous):

^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

"729 3^9 = 19,683 (3*3*3*3*3*3*3*3*3) 3^-3 = 1/27 (1/3 * 3 *3, aka 1/27) Then multiply 19,683 by 1/27 Which would be 729." -Racecar9872

OpenStudy (igreen):

@racecar9872 Please no answers.

OpenStudy (anonymous):

@iGreen sorry I'm new to this site